OpenStudy (anonymous):

a 4 kg. block is accelerated upward by a cord whose breaking strength is 60 newton. find the maximum acceleration which can be given the block without breaking the cord.please help me.

OpenStudy (souvik):

 F-Mg=Ma a=F/M-g here F=60N M=4 kg g=9.8 m/s^2 ~ 10 m/s^2

OpenStudy (anonymous):

thank you!